# Beyond the Trade-off Curve: Multivariate and Advanced Risk-Utility Maps for Evaluating Anonymized and Synthetic Data

### v2

@@ Abstract
Anonymizing microdata requires balancing the reduction of disclosure risk with the preservation of data utility. Traditional evaluations often rely on single measures or two-dimensional risk-utility (R-U) maps, but real-world assessments involve multiple, often correlated, indicators of both risk and utility. Pairwise comparisons of these measures can be inefficient and incomplete. We therefore systematically compare six visualization approaches for simultaneous evaluation of multiple risk and utility measures: heatmaps, dot plots, composite scatterplots, parallel coordinate plots, radial profile charts, and PCA-based biplots. We introduce blockwise PCA for composite scatterplots and joint PCA for biplots that simultaneously reveal method performance and measure interrelationships. Through systematic identification of Pareto-optimal methods in all approaches, we demonstrate how multivariate visualization supports a more informed selection of anonymization methods.
Keywords: statistical disclosure control, risk-utility, Pareto optimality, multivariate statistics, visualization
@@ 1 Introduction
In microdata anonymization, modifications to the original dataset, such as suppression, generalization, perturbation, or synthetic data generation, are essential to reduce disclosure risk [21, 47]. However, these modifications inevitably result in a loss of information, potentially limiting the analytical value of the data. Effective anonymization therefore requires balancing the confidentiality gained from lowering disclosure risk with the preservation of data utility. A possibility for evaluating this trade-off visually is the risk–utility (R-U) confidentiality map [8]. R–U maps depict anonymized datasets according to their measured disclosure risk and data utility, facilitating the systematic comparison of alternative anonymization approaches. By anonymization approaches, we refer to the different strategies by which datasets can be anonymized. These may involve method-specific variations, such as adjusting the level of noise in noise addition [1] or comparisons between methods – e.g., the effect of microaggregation [6] and post randomization [13] on disclosure risk. They can also include differences in synthetic data, either through employing distinct data generators and/or by producing multiple datasets with the same generator.
@@
In practice, it is common to report one or more disclosure risk indicators and utility measures separately, often presented as summary statistics in tables or text, or visualized through simplified two-dimensional R–U maps [31, 48, 17, 24, 23]. Although such representations, especially R-U maps, aid in interpretation, displaying several risk and utility quickly becomes cumbersome. One remedy is to use faceting and arrange multiple R–U maps as small multiples: by enforcing comparisons of change, of differences between objects, small multiples are often the best solution for a wide range of presentation problems [52, p. 67-68]. However, as the number of measures grows and thereby the number of panels needed for their display, this approach becomes more fragmented and inefficient [18]. For example, with five distinct risk measures and five utility measures, 25 different R–U plots would be required to examine all pairwise relationships, complicating overall interpretation.
@@
This challenge stems from the inherently multivariate nature of the evaluation problem: different risk and utility measures reflect distinct, often uncorrelated, aspects of data protection and analytical validity. Presenting them separately obscures potential interactions and trade-offs across dimensions. A comprehensive assessment therefore requires methods that jointly account for this multivariate structure, moving beyond simple pairwise comparisons to enable simultaneous consideration of multiple criteria. By framing risk and utility as a multi-objective optimization problem, we provide tools for more holistic evaluation of data releases – whether synthetically or traditionally anonymized [5, 4].
@@
In this article, we extend the classical risk-utility map to a multivariate setting through systematic comparison of six visualization approaches: heatmaps, dot plots, composite scatterplots, parallel coordinate plots, radial profile charts, and PCA-based biplots. We introduce three methodological innovations using principal component analysis: (1) blockwise PCA that extracts principal components separately for risk and utility measure blocks, (2) alignment analysis that validates dimensionality reduction by correlating composite measures with principal components, and (3) PCA-based biplots that simultaneously visualize method performances and measure relationships. To our knowledge, this represents the first systematic application of PCA to risk-utility visualization in statistical disclosure control. We further demonstrate systematic Pareto-optimal approach identification across all visualization approaches and evaluate each approach across twelve analytical capabilities. The design of these tools follows Tufte’s data-ink principle [51] and key guidelines for visual data communication [10].
The remainder of the paper is structured as follows. Section 2 introduces the conceptual and methodological foundation, framing evaluation as multi-objective optimization and presenting our visualization toolkit. Section 3 applies the framework to synthetic microdata from the EU-SILC dataset, illustrating how different approaches reveal performance trade-offs among synthetic data generators (SDGs). Section 4 discusses practical implications, method-specific considerations, and future directions. Section 5 provides key recommendations for practitioners and synthesizes our main contributions.
2 Methodology
This section develops a visual approach to the multi-objective optimization problem underlying the disclosure risk-utility trade-off. We introduce methods well suited to evaluating and visualizing anonymized and synthesized data across multiple risk and utility dimensions.
@@ 2.1 Pareto-Optimal/Efficient Trade-offs
The multivariate evaluation of disclosure risk and data utility naturally constitutes a multi-objective optimization problem, since improvements in one dimension often come at the expense of another. In such settings, the objectives are said to be at least partly conflicting [29, p. 5], and a solution (or observation) is called non-dominated, or Pareto-optimal/Pareto-efficient, if none of the objectives can be improved without degrading at least one of the others. Intuitively, this means that a Pareto-optimal solution represents an efficient trade-off: it cannot be improved in one aspect (e.g., utility) without performing worse in another (e.g., risk). Any method that is not Pareto-optimal is strictly inferior to at least one alternative [25, p. 547], a concept originating with Vilfredo Pareto’s Cours d’économie politique (1896–1897).
@@
with at least one strict inequality [29, p. 11] [9, Def. 5, p. 588]. An anonymization approach $i$ is Pareto-optimal if there exists no other anonymization approach $j$ that dominates it. This corresponds to strong Pareto optimality, since at least one inequality must be strict; the weaker notion, which also allows ties, is not considered here [29, p. 19].
@@
Additional decision criteria can also be applied. One option is to impose a risk-tolerance threshold and select the Pareto-optimal solution with the highest utility subject to this bound. Another intuitive “bang-for-buck” heuristic, applicable when the frontier is smooth and concave toward the origin, is to choose the knee (or elbow) point – where marginal increases in risk begin to yield only diminishing gains in utility, corresponding to the location of greatest curvature along the curve [50, p. 275].
In practice, Pareto-optimal observations can be highlighted in visualizations using different colors or shapes, thereby conveying multidimensional information within a two-dimensional plot [32]. However, the frontier itself can only be directly visualized in two dimensions.
2.2 Visualization and Evaluation Tools
@@
As argued in Section 1, the risk-utility problem is fundamentally multivariate – joint feasibility matters more than marginal checks. We therefore introduce concise visualization strategies: different PCA biplots, R-U maps, origami charts and related multivariate views with Pareto-optimal solutions highlighted. We illustrate these on real data in Section 3.
2.2.1 Heatmaps

### v3

# Beyond the Trade-off Curve: Multivariate and Advanced Risk-Utility Maps for Evaluating Anonymized Data

@@ Abstract
Anonymizing microdata requires balancing disclosure risk reduction with the preservation of data utility. Traditional evaluations often rely on single measures or two-dimensional risk-utility (R-U) maps, but real-world assessments involve multiple, often correlated, indicators of both risk and utility — a fundamentally multivariate problem that pairwise comparisons fail to capture both efficiently and completely. We systematically compare six visualization approaches for simultaneous evaluation of multiple risk and utility measures: heatmaps, dot plots, composite scatterplots, parallel coordinate plots, radial profile charts, and principal component analysis (PCA)-based biplots. We introduce blockwise PCA for composite scatterplots and joint PCA for biplots that simultaneously reveal method performance and measure interrelationships, and apply systematic Pareto-optimal method identification across all approaches where applicable, with dominance assessed in the original composite score space. Our comparison shows that no single approach dominates across all criteria: PCA biplots best reveal multivariate structure, while composite scatterplots offer intuitive summaries accessible to broader audiences. Combining complementary visualizations provides the most complete basis for evaluating the risk-utility trade-off.
Keywords: statistical disclosure control, risk-utility, Pareto optimality, multivariate statistics, visualization, RU-map, synthetic data
@@ 1 Introduction
In microdata anonymization, modifications to the original dataset, such as suppression, generalization, perturbation, or synthetic data generation, are essential to reduce disclosure risk [2012_Hundepool, 2017_Templ]. However, these modifications inevitably result in a loss of information, potentially limiting the analytical value of the data. Effective anonymization therefore requires balancing the confidentiality gained from lowering disclosure risk with the preservation of data utility. A possibility for evaluating this trade-off visually is the risk–utility (R-U) confidentiality map [2001_Duncan]. R–U maps depict anonymized datasets according to their measured disclosure risk and data utility, facilitating the systematic comparison of alternative anonymization approaches. By anonymization approaches, we refer to the different strategies by which datasets can be anonymized. These may involve method-specific variations, such as adjusting the level of noise in noise addition [Brand_2002_Microdata] or comparisons between methods – e.g., the effect of microaggregation [Defays_1998_Masking] and post randomization [Gouweleeuw_1998_Post] on disclosure risk. They can also include differences in synthetic data, either through employing distinct data generators and/or by producing multiple datasets with the same generator. Although the visualization methods presented in this paper are illustrated using synthetic data examples, the proposed visualizations can also be applied to traditional anonymization methods. The prerequisite is that the utility and risk measures used for comparison are appropriate and comparable across all applied anonymization approaches.
@@
In practice, it is common to report one or more disclosure risk indicators and utility measures separately, often presented as summary statistics in tables or text, or visualized through simplified two-dimensional R–U maps [[, e.g.,]]Muralidhar_2006_Data, Templ_2008_Robust, Hornby_2021_Identification, Little_2022_Comparing, Little_2025_Synthetica. Although such representations, especially R-U maps, aid in interpretation, displaying several risk and utility measures quickly becomes cumbersome. One remedy is to use faceting and arrange multiple R–U maps as small multiples: by enforcing comparisons of change, of differences between objects, small multiples are often the best solution for a wide range of presentation problems [Tufte_1990_Envisioning, p. 67-68]. However, as the number of measures grows and thereby the number of panels needed for their display, this approach becomes more fragmented and inefficient [Hosseinpour_2025_Examining]. For example, with five distinct risk measures and five utility measures, 25 different R–U plots would be required to examine all pairwise relationships, complicating overall interpretation.
@@
This challenge stems from the inherently multivariate nature of the evaluation problem: different risk and utility measures reflect distinct, often uncorrelated, aspects of data protection and analytical validity. Presenting them separately obscures potential interactions and trade-offs across dimensions. While [2022_Dankar_Multi] explicitly acknowledge the multidimensional nature of utility by highlighting the large number of available utility metrics and the absence of general guidelines or standardised thresholds for their interpretation, they also propose dimension-reduction approaches, such as principal component analysis, to summarize multiple utility measures [2022_Dankar_PCA]. However, comparable considerations jointly spanning both risk and utility dimensions remain limited. A comprehensive assessment therefore requires methods that jointly account for this multivariate structure, moving beyond simple pairwise comparisons to enable simultaneous consideration of multiple criteria. By framing risk and utility as a multi-objective optimization problem, we provide tools for more holistic evaluation of data releases – whether synthetically or traditionally anonymized.
@@
In this article, we extend the classical risk-utility map to a multivariate setting through systematic comparison of six visualization approaches: heatmaps, dot plots, composite scatterplots, parallel coordinate plots, radial profile charts, and PCA-based biplots. We introduce three methodological innovations using principal component analysis: (1) blockwise PCA that extracts principal components separately for risk and utility measure blocks, (2) alignment analysis that validates dimensionality reduction by correlating composite measures with principal components, and (3) PCA-based biplots that simultaneously visualize method performances and measure relationships. To our knowledge, this represents the first systematic application of PCA to risk-utility visualization in statistical disclosure control. We further demonstrate systematic Pareto-optimal approach identification across all visualization approaches and evaluate each approach across twelve evaluation criteria. The design of these tools follows Tufte’s data-ink principle [Tufte_1983_Visual] and key guidelines for visual data communication [Franconeri_2021_Science].
The remainder of the paper is organized as follows. Section 2 introduces Pareto optimality and discusses key considerations related to the scaling of risk and utility measures. Section 3 describes the example data used for illustration, presents the visualization methods, and discusses their applications; a systematic comparison of the methods is provided in Section 4. Section 5 explores practical implications and outlines directions for future research. Finally, Section 6 summarizes the main contributions of the paper and offers key recommendations for practitioners. The visualization methods are presented in order of increasing analytical depth, from tabular overviews to dimensionality-reduction-based approaches.
2 Methodological Framework
@@ 2.1 Pareto-Optimal/Efficient Trade-offs
The multivariate evaluation of disclosure risk and data utility naturally constitutes a multi-objective optimization problem, since improvements in one dimension often come at the expense of another. In such settings, the objectives are said to be at least partly conflicting [Miettinen_1998_Nonlinear, p. 5], and a solution (or observation) is called non-dominated, or Pareto-optimal/Pareto-efficient, if none of the objectives can be improved without degrading at least one of the others. Intuitively, this means that a Pareto-optimal solution represents an efficient trade-off: it cannot be improved in one aspect (e.g., utility) without performing worse in another (e.g., risk). Any method that is not Pareto-optimal is strictly inferior to at least one alternative [MasColell_1995_Chapter, p. 547], a concept originating with Vilfredo Pareto’s Cours d’économie politique (1896–1897).
@@
with at least one strict inequality [Miettinen_1998_Nonlinear, p. 11] [Emmerich_2018_Tutorial, Def. 5, p. 588]. An anonymization approach $i$ is Pareto-optimal if there exists no other anonymization approach $j$ that dominates it. This corresponds to strong Pareto optimality, since at least one inequality must be strict; the weaker notion, which also allows ties, is not considered here [Miettinen_1998_Nonlinear, p. 19].
@@
Additional decision criteria can also be applied. One option is to impose a risk-tolerance threshold and select the Pareto-optimal solution with the highest utility subject to this bound. Another intuitive “bang-for-buck” heuristic, applicable when the frontier is smooth and concave toward the origin, is to choose the knee (or elbow) point – where marginal increases in risk begin to yield only diminishing gains in utility, corresponding to the location of greatest curvature along the curve [Thorndike_1953_Who, p. 275].
In practice, Pareto-optimal observations can be highlighted in visualizations using different colors or shapes [Nagar_2023_Visualization], but the frontier itself can only be directly visualized in two dimensions. Pareto-optimality can be evaluated either on the full vector of raw measures or on aggregated composite scores. The former is theoretically cleaner – strictly monotone transformations preserve dominance metric-by-metric – but inconsistent with two-dimensional visualization and prone to degeneracy: with many measures, it becomes increasingly difficult for any approach to dominate another on all dimensions simultaneously, causing most approaches to appear non-dominated. The latter reduces the problem to two dimensions at the cost of making the identified set sensitive to aggregation and scaling choices.
2.2 Orientation and Scaling of Risk and Utility Measures
Since risk and utility measures capture fundamentally different aspects of data properties (e.g., disclosure probabilities and distance distributions), they are naturally defined on different scales, units, and directions. Whether higher values indicate better or worse outcomes depends on the measure’s definition: for instance, disclosure risk may quantify the attacker’s success (higher = worse) or the defender’s success (higher = better), and distributional utility may be expressed as distance (lower = better) or as similarity (higher = better). Moreover, measures differ in range and units – some are bounded probabilities on $[0,1]$, others are unbounded distances – making direct comparison across measures impossible without rescaling. For joint multivariate analysis and visualization, both a consistent orientation and a transformation to a common scale are therefore necessary.
When measures within a category have mixed orientations (e.g., some utility measures are distance-based where lower = better, while others are similarity-based where higher = better), we recommend reorienting each individually before scaling. Whether risk and utility should share a common direction depends on the intended visualization: separate orientations are natural for dual-panel displays such as dot plots (Section 3.2), split parallel coordinate plots (PCPs; Section 3.4) and heatmaps (Section 3.1) with direction-encoded color scales, while a unified direction is recommended for combined radar charts (Section 3.5), single-panel PCPs, and threshold-based scaling approaches where a single threshold should carry the same interpretation across all axes. When all measures within each category already share the same direction, a single reversal, $x^{\prime}=1-x$ after min–max scaling, or $x^{\prime}=-x$ after z-score standardization, can reorient an entire category (i.e., utility or risk) for intuitive visual interpretation.
Multivariate visualizations also require comparable axis ranges. Radar charts need bounded scales to produce meaningful polygon shapes, parallel coordinate plots benefit from comparable ranges for visual readability, and PCA biplots require standardization to prevent variables with larger ranges from dominating the principal components [Greenacre_2010_Biplots]. The choice of scaling method depends on the analytical objective, and conventional normalization approaches do not always yield high-quality visual or structural representations [Dierkes_2025_Scalingevariant]. Min–max scaling to $[0,1]$ is particularly suitable for radar charts and parallel coordinate plots, as 0 and 1 correspond to the observed worst and best performance across synthetic data generators (SDGs), directly conveying the attainable performance range. z-score standardization is generally preferred for PCA biplots, as it equalizes variances across variables. However, z-score values express performance relative to the mean in standard deviation units, which is less intuitive for visualizing performance trade-offs than the bounded $[0,1]$ range of min–max scaling.
Both of these transformations necessarily abstract from the original metric units, which implies that absolute interpretability of thresholds – such as maximum acceptable disclosure risk levels – cannot be directly represented in the scaled space. Alternative approaches that explicitly incorporate such acceptability criteria include desirability functions [Derringer_1980_Simultaneous], which map raw values to a [0,1] scale using threshold-anchored transformations, and sigmoid functions, which provide smooth bounded mappings centered around a target value. While some measures have well-established baselines, thresholds for many synthetic data quality measures are context-dependent and must be defined for the specific use case and data environment [Drechsler_2022_Challenges]. Table 1 compares these approaches across four desirable scaling properties – boundedness, a universal threshold, full differentiation, and fixed variance – illustrating that no single approach satisfies all criteria simultaneously.
Table 1: Desirable scaling properties
[…]
Min–max $\frac{x-x_{\min}}{x_{\max}-x_{\min}}$ $\times$ $\times$
z-score $\frac{x-\bar{x}}{s}$ $\times$ $\times$
Threshold-proportional $\frac{x}{\tau}$ $\times$ $\times$
Sigmoid $\frac{1}{1+e^{-k(x-\tau)}}$ $\sim*$ $\times$
Desirability $d_{i}(x_{i})$ $\times**$ $\times$
  • $\tau$: acceptability threshold; $k$: steepness parameter; $d_{i}$: individual desirability function.
    Soft approximation; loses differentiation in the tails.
    Differentiates within the acceptable range; assigns zero outside it.
  • $\dagger$
    Requires consistent orientation of all measures so that $\tau$ carries a uniform interpretation across all axes.
When Pareto dominance is evaluated directly on the full vector of risk and utility measures, strictly monotone transformations of individual metrics do not alter the set of non-dominated solutions, since dominance is defined through metric-by-metric comparisons and strictly monotone transformations preserve the ordering within each metric. However, the orientation of each measure must be consistent with the optimization direction. When dominance is instead assessed on aggregated composite scores (e.g., mean risk and mean utility), scaling implicitly determines the relative contribution of each measure to the composite, and consequently the identified Pareto set may differ depending on the chosen scaling transformation (see Appendix B for a numerical illustration).
3 Visualization and Evaluation Tools
@@
This section develops a visual approach to the multi-objective optimization problem underlying the disclosure risk-utility trade-off. We introduce methods well suited to evaluating and visualizing anonymized and synthesized data across multiple risk and utility dimensions. As argued in Section 1, the risk-utility problem is fundamentally multivariate. We therefore introduce multivariate visualization strategies: different PCA biplots, R-U maps, origami plots and related multivariate views with Pareto-optimal solutions highlighted.
To demonstrate the proposed visualization methods, data from the European Union Statistics on Income and Living Conditions (EU-SILC) are utilized. EU-SILC is a flagship data source in official statistics, widely used to monitor poverty, social inclusion, and related policy targets in Europe. We use the Austrian EU-SILC public-use file from 2013. The data has a hierarchical structure, with individuals nested within households. A concise description of key variables and further details is available in the manual of the R package simPop [Templ_2017_Simulation]; a comprehensive variable catalogue is provided in [silc] and online at https://www.gesis.org/en/missy/materials/EU-SILC/documents/codebooks (accessed 2025-06-25).
We synthesize data using a methodologically diverse set of commonly used synthetic data generators (SDGs), each applied 10 times to account for stochastic variation. The SDGs and their software packages are described in Table 6 in Appendix A; SDG-specific parameter settings used for the EU-SILC synthesis are provided in Table 8 of the same appendix.
Over the past 30 years, the field of synthetic data has grown substantially, giving rise to a large number of evaluation metrics [Kaabachi_2025_Scoping]. However, no clear consensus has emerged on a standard set of measures [Drechsler_2024_30]. We therefore select a diverse range of metrics spanning different categories of both utility and risk. For utility, we draw on the taxonomy of [Drechsler_2024_30] and choose among global utility measures (e.g. pMSE), outcome-specific utility measures (e.g. Confidence Interval Proximity), and fit-for-purpose measures (e.g. invalid Households with only children). For risk, we draw on the recently published consensus study on privacy metrics for synthetic data [Pilgram_2025_Consensus], covering three disclosure categories: identity disclosure (RepU, DCR), attribute disclosure (TCAP, RAPID), and membership inference (MIA). These measures and their descriptions are provided in Table 2 (for further details, see Appendix A, Table 7).The MIA implementation is based on the SynthEval framework [Lautrup_2025_Syntheval].
We emphasize that the chosen SDGs and evaluation metrics are intended solely to support the proposed visualizations and should not be interpreted as a systematic comparison or benchmarking of different SDGs. The selection of both generators and metrics was made subjectively and is not central to the contribution of this paper, which focuses on proposing and evaluating visualizations for the multivariate nature of the risk-utility trade-off. Accordingly, the SDGs were not extensively tuned to the specific dataset, and the results should not be interpreted as performance comparisons between SDGs.
Table 2: Description of the risk and utility measures.
Type Measure & Abbreviation Author / Used in Description Risk Replicated Uniques (RepU) [Raab_2025_Practical, Raab_2025_Confidentiality, e.g.,] Percentage of replicated sample uniques in synthetic dataset. Risk Distance to Closest Record (DCR) [Yao_2025_DCR, e.g.,] Ratio of mean nearest-neighbour distances from synthetic records to training data versus holdout. Risk Membership Inference Attack (MIA) [Lautrup_2025_Syntheval, ElEmam_2022_Validating, Houssiau_2022_TAPAS, Shokri_2017_Membership, e.g.,] Probability that an attacker can correctly infer whether a record was part of the training data. Risk Targeted Correct Attribution Probability (TCAP) [Taub_2019_Creating] Probability of correct attribute inference. Risk Risk of Attribute Prediction-Induced Disclosure (RAPID) [Templ_2026_RAPID] Expected share of predicted attribute disclosure. Utility Confidence Interval Proximity (CIProx) [Karr_2006_Frameworka, Drechsler_2009_Disclosure, e.g.,] Deviation of confidence intervals for a sensitive variable. Utility Propensity Mean Squared Error (pMSE) [snoke_general_2017, Raab_2021_Assessing, e.g.,] Predictive score from distinguishing real vs. synthetic data. Utility Wasserstein Distance (Wasserstein) [Vasershteyn_1969_Markovskie] Wasserstein distance between numeric distributions. Utility Households with only children (NoAdultHH) [thees24, e.g.,] Households in which all members are under 18 years of age. Since such households are demographically impossible in the EU-SILC data structure, the count should be zero in valid synthetic data. A non-zero value indicates a logical consistency violation. Utility Correlation Matrices Differences (CMD) [Miletic_2025_Utilitybased, e.g.,] Mean absolute difference between corresponding real and synthetic correlation coefficients. Note: CIProx adapts the confidence interval overlap measure, applying it to variable-level means rather than regression estimands.
Regarding orientation, all utility measures are aligned such that higher values indicate better performance. For risk measures, lower values indicate lower disclosure risk; an exception is the Distance to Closest Record (DCR), which by construction yields higher values for lower risk and was therefore inverted to ensure a consistent orientation across all risk measures.
This orientation is maintained throughout the analysis. For certain visualizations that require a unified interpretation of radial extent (e.g., origami plots), risk measures are temporarily inverted so that larger values consistently correspond to better performance across all axes; this transformation is applied solely for visual interpretability and does not affect any quantitative analysis.
For scaling, different transformations are used depending on the analytical objective. Min–max scaling to the interval [0,1] is applied for composite R–U maps and bounded multivariate visualizations such as heatmaps, dot plots, and radial charts, where interpretability of the observed performance range is desirable. Scaling is performed per metric across all SDGs.
Pareto-optimality is assessed on the min–max scaled composite scores – mean utility and mean risk. Although min–max scaling is sensitive to extreme values, the set of SDGs in our setting is fixed, and extreme observations – including the original dataset as a reference point – represent meaningful benchmark anchors rather than nuisance outliers. For PCA-based visualizations, z-score standardization is applied to ensure comparability across measures.
To verify this choice, we compared PCA biplots under min–max scaling, z-score standardization, and the scaling optimization algorithm proposed by [Dierkes_2025_Scalingevariant], which uses Nelder–Mead optimization to determine per-dimension scaling factors that maximize a visual quality criterion for two-dimensional projections. Min–max and z-score scaling produced broadly similar projections with comparable cluster structure and synthesizer separation, confirming that the visualization is robust to the choice between these two common scaling approaches for the present data. The Dierkes optimization, applied using the Data Space Ratio as the unsupervised quality criterion, collapsed the projection to a near-one-dimensional solution and was therefore discarded. We proceed with min–max scaling for bounded visualizations and z-score standardization for PCA, as described above. All risk and utility measure calculations, data manipulation, and visualization were performed in R [RCoreTeam_2025_Language] [^1] We mainly used the tidyverse ecosystem [Wickham_2019_Welcome], specifically ggplot2 [Wickham_2016_Ggplot2] for visualization. Additional packages are cited where used..
3.1 Heatmaps